# Structured Multi-Track Accompaniment Arrangement via Style Prior Modelling

### v3

@@ 1 Introduction
Representation learning techniques have enabled new possibilities for controllable generative modelling. By learning implicit style representations, which are often hard to explicitly label (e.g., timbre of music audio [21], texture of music composition [39], and artistic style in paintings [20]), new music and artworks can be created via style transfer and latent space sampling. These learned style factors can also serve as external controls for downstream generative models, including Transformers [18, 36] and diffusion models [42]. HowNever, applying style factors to long-term sequence generation remains a challenging task. Existing approaches rely on style templates specified manually or by heuristic rules [36, 42, 51], which are impractical for long-term generation. Moreover, when structural constraints are imposed, misaligned style factors can result in incoherent outputs.
@@
In this paper, we study style prior modelling through the task of multi-track accompaniment arrangement, a typical scenario for long-term conditional sequence generation. We assume the input of a piano accompaniment score, which typically carries a verse-chorus structure. Our target is to generate corresponding multi-track arrangements featuring band orchestration. We start by disentangling a band score at time $t$ into piano reduction $\mathbf{c}_{t}$ (content factor) and orchestral function $\mathbf{s}_{t}^{k}$ (style factors for individual tracks $k=1,2,\cdots,K$). On top of this, we model the prior of finding appropriate functions to orchestrate a given piano score, or formally $p(\mathbf{s}_{1:T}^{1:K}\mid\mathbf{c}_{1:T})$. To model dependencies on both time ($T$) and track ($K$) directions, we develop a multi-stream Transformer with interleaved time-wise and track-wise layers. The track-wise layer allows for flexible control over the choice of instruments and the number of tracks, while the time-wise layer ensures structural alignment through cross-attention to the piano reduction. Decoding the inferred $\mathbf{s}_{1:T}^{1:K}$ with $\mathbf{c}_{1:T}$, we can address accompaniment arrangement in a flexible multi-track form with extended whole-song structure.
@@ 3.1 Data Representation
Latent Dimension $\mathbf{z}\in\mathbb{R}^{T\times 256}$ $\mathbf{c}\in\mathbb{R}^{T\times 256}$ $s\in[0..127]^{8T\times K}$
@@ 4 Whole-Song Multi-Track Accompaniment Arrangement
We showcase an arrangement example by the complete system in Figure 4. The system input is a lead sheet shown by the $\tt{Mel}$ staff. The final output is the accompaniment in the rest staves. Notably, the lead sheet consists of 60 bars in an structure of $\tt{i4A8B8B8x4A8B8B8O4}$ (using notations by [5]). Here, $\tt{i4}$, $\tt{x4}$, and $\tt{O4}$ each denote a 4-bar intro, interlude, and outro. $\tt{A8}$ and $\tt{B8}$ represent an 8-bar verse and chorus, respectively. Figure 4 shows the arrangement result for the first and third choruses, spanning from bar 13 to 41. We leverage control option 2 to customize the instrumentation as celesta, acoustic guitars (2), electric pianos (2), acoustic piano, violin, brass, and electric bass in a total of $K=9$ tracks. The complete arrangement score is included in Appendix E.
@@ 5.3 Objective Evaluation on Multi-Track Arrangement
Ours $\mathbf{0.564}\pm 0.014^{a}$ $\mathbf{0.300}\pm 0.005^{a}$ $\mathbf{1.519}\pm 0.030^{a}$ $0.461\pm 0.005^{b}$
@@ 5.6 Ablation Study on Piano Arrangement
Table 4: Ablation study on alternative lead sheet to piano arrangement (i.e., Stage 1) modules. Here we investigate the impact of Stage 1 to the entire two-stage system. Evaluation results are based on the final multi-track arrangement using respective Stage 1 modules.
@@
Ours (Stage 1 + Stage 2) $\mathbf{0.564}\pm 0.014$ $\mathbf{1.519}\pm 0.030$ $\mathbf{0.300}\pm 0.005$
@@
Piano Arr. Module Chord Acc. $\uparrow$ Structure $\uparrow$
@@ A.1 Autoencoder
The autoencoder consists of a VQ-VAE submodule and an overarching VAE. The encoder of the VQ-VAE consists of a 1-D convolutional layer of kernal size 4, stride 4, and 16 output channels, followed by a vector quantization block with codebook size 128. The decoder takes the concatenated latent codes and leverages two fully-connected layers (shape $128\times 256$ and $256\times 32$) for reconstruction. In the overarching VAE, Piano Encoder and Track Decoder are adapted from PianoTree VAE [40]. The encoder first applies a pitch-wise bi-directional GRU to summarize concurrent notes at time step $n$ and then applies a time-wise GRU to encode the full representation. The decoder mirrors the encoder structure with time- and pitch-wise uni-directional GRUs to reconstruct individual tracks. We use hidden size 256 in a single layer for pitch GRUs and 512 for time GRUs. The Track Separator is a 2-layer Transformer encoder with 8 attention heads, 0.1 dropout ratio, and GELU activation [12]. The hidden dimensions of self-attention $d_{\mathrm{model}}$ and feed-forward layers $d_{\mathrm{ff}}$ are 512 and 1024, respectively.
@@ B.1 Degree of Arrangement
where $\overline{d}$ is the mean. $K$ is the total number of tracks. To establish a reference point, we calculate the ground-truth $\mathrm{DOA}=0.333$ based on 1000 randomly selected pieces from the LMD dataset. Within this context, a higher $\mathrm{DOA}$ signifies a more creative arrangement.
@@ B.2 Strcture Awareness
where ||$\cdot$|| ⋅ | is the cardinality of a set. $T$ is the number of 2-bar segments. When applying $\mathrm{ILS}$, we use [5] to automatically lable the phrase structure of a piece. To establish a reference point, we calculate the ground-truth $\mathrm{ILS}=1.980$ based on the POP909 dataset (with phrase annotation by human). Within this context, a higher $\mathrm{ILS}$ signifies saliency with long-term phrase-level structure.
@@ Appendix C Experiment on Noise Weight $\gamma$
Continuing from Section 3.3, we compare different $\gamma$ values and see their impact to the model performance. When applying our model to piano to multi-track arrangement, $\gamma$ balances the force of a noisy factor added to the piano, which encourages a partial unconditional generation capability. The experimental settings are the same as Section 5.5. We evaluate the results based on faithfulness and DOA. In Table 6, we report mean value, standard error of mean ($\mathrm{sem}$), and statistical significance computed by Wilcoxon signed rank test. By varying the $\gamma$ value, we observe a controllable balance between faithfulness and creativity. Specifically, a larger $\gamma$ encourages creativity (higher DOA) at the cost of faithfulness. If not mentioned otherwise, we use $\gamma=0.25$ for experiments in this paper.
@@ Appendix G Broader Impacts
However, we acknowledge the need to address potential risks. The accessibility of our system may inadvertently lead to excessive reliance on automation, potentially impeding the development of fundamental skills among musicians. Additionally, widespread adoption of the system may contribute to the homogenization of music, threatening the distinctiveness and individuality that are crucial to artistic expression. We recognize that our datasets predominantly features contemporary Western music, which introduces a cultural bias that could limit the diversity of generated compositions.

### v4

@@ 1 Introduction
Representation learning techniques have enabled new possibilities for controllable generative modelling. By learning implicit style representations, which are often hard to explicitly label (e.g., timbre of music audio [21], texture of music composition [39], and artistic style in paintings [20]), new music and artworks can be created via style transfer and latent space sampling. These learned style factors can also serve as external controls for downstream generative models, including Transformers [18, 36] and diffusion models [42]. However, applying style factors to long-term sequence generation remains a challenging task. Existing approaches rely on style templates specified manually or by heuristic rules [36, 42, 51], which are impractical for long-term generation. Moreover, when structural constraints are imposed, misaligned style factors can result in incoherent outputs.
@@
In this paper, we study style prior modelling through the task of multi-track accompaniment arrangement, a typical scenario for long-term conditional sequence generation. We assume the input of a piano accompaniment score, which typically carries a verse-chorus structure. Our target is to generate corresponding multi-track arrangements featuring band orchestration. We start by disentangling a band score at time $t$ into piano reduction $\mathbf{c}_{t}$ (content factor) and orchestral function $\mathbf{s}_{t}^{k}$ (style factors for individual tracks $k=1,2,\cdots,K$). On top of this, we model the prior of finding appropriate functions to orchestrate a given piano score, or formally $p(\mathbf{s}_{1:T}^{1:K}\mid\mathbf{c}_{1:T})$. To model dependencies in both time ($T$) and track ($K$) directions, we develop a multi-stream Transformer with interleaved time-wise and track-wise layers. The track-wise layer allows for flexible control over the choice of instruments and the number of tracks, while the time-wise layer ensures structural alignment through cross-attention to the piano reduction. Decoding the inferred $\mathbf{s}_{1:T}^{1:K}$ with $\mathbf{c}_{1:T}$, we can address accompaniment arrangement in a flexible multi-track form with extended whole-song structure.
@@ 3.1 Data Representation
Latent Dimension $\mathbf{z}\in\mathbb{R}^{T\times K\times 256}$ $\mathbf{c}\in\mathbb{R}^{T\times 256}$ $s\in[0..127]^{8T\times K}$
@@ 4 Whole-Song Multi-Track Accompaniment Arrangement
We showcase an arrangement example by the complete system in Figure 4. The system input is a lead sheet shown by the $\tt{Mel}$ staff. The final output is the accompaniment in the rest staves. Notably, the lead sheet consists of 60 bars in a structure of $\tt{i4A8B8B8x4A8B8B8O4}$ (using notations by [5]). Here, $\tt{i4}$, $\tt{x4}$, and $\tt{O4}$ each denote a 4-bar intro, interlude, and outro. $\tt{A8}$ and $\tt{B8}$ represent an 8-bar verse and chorus, respectively. Figure 4 shows the arrangement result for the first and third choruses, spanning from bar 13 to 41. We leverage control option 2 to customize the instrumentation as celesta, acoustic guitars (2), electric pianos (2), acoustic piano, violin, brass, and electric bass in a total of $K=9$ tracks. The complete arrangement score is included in Appendix E.
@@ 5.3 Objective Evaluation on Multi-Track Arrangement
Ours $\mathbf{0.567}\pm 0.014^{a}$ $\mathbf{0.300}\pm 0.004^{a}$ $\mathbf{1.520}\pm 0.030^{a}$ $0.461\pm 0.005^{b}$
@@ 5.6 Ablation Study on Piano Arrangement
Table 4: Ablation study on alternative lead sheet to piano arrangement (i.e., Stage 1) modules. Here we investigate the impact of Stage 1 on the entire two-stage system. Evaluation results are based on the final multi-track arrangement using respective Stage 1 modules.
@@
Ours (Stage 1 + Stage 2) $\mathbf{0.567}\pm 0.014$ $\mathbf{1.520}\pm 0.030$ $\mathbf{0.300}\pm 0.004$
@@
Piano Arrangement Module Chord Acc. $\uparrow$ Structure $\uparrow$
@@ A.1 Autoencoder
The autoencoder consists of a VQ-VAE submodule and an overarching VAE. The encoder of the VQ-VAE consists of a 1-D convolutional layer of kernel size 4, stride 4, and 16 output channels, followed by a vector quantization block with codebook size 128. The decoder takes the concatenated latent codes and leverages two fully-connected layers (shape $128\times 256$ and $256\times 32$) for reconstruction. In the overarching VAE, Piano Encoder and Track Decoder are adapted from PianoTree VAE [40]. The encoder first applies a pitch-wise bi-directional GRU to summarize concurrent notes at time step $n$ and then applies a time-wise GRU to encode the full representation. The decoder mirrors the encoder structure with time- and pitch-wise uni-directional GRUs to reconstruct individual tracks. We use hidden size 256 in a single layer for pitch GRUs and 512 for time GRUs. The Track Separator is a 2-layer Transformer encoder with 8 attention heads, 0.1 dropout ratio, and GELU activation [12]. The hidden dimensions of self-attention $d_{\mathrm{model}}$ and feed-forward layers $d_{\mathrm{ff}}$ are 512 and 1024, respectively.
@@ B.1 Degree of Arrangement
where $\overline{d}$ is the mean. $K$ is the total number of tracks. To establish a reference point, we calculate the ground-truth $\mathrm{DOA}=0.333$ based on 1000 random pieces from the LMD dataset with at least 5 tracks. Within this context, a higher $\mathrm{DOA}$ signifies a more creative arrangement.
@@ B.2 Strcture Awareness
where ||$\cdot$|| ⋅ | is the cardinality of a set. $T$ is the number of 2-bar segments. When applying $\mathrm{ILS}$, we use [5] to automatically label the phrase structure of a piece. To establish a reference point, we calculate the ground-truth $\mathrm{ILS}=1.980$ based on the POP909 dataset (with phrase annotation by human). Within this context, a higher $\mathrm{ILS}$ signifies saliency with long-term phrase-level structure.
@@ Appendix C Experiment on Noise Weight $\gamma$
Continuing from Section 3.3, we compare different $\gamma$ values and see their impact on the model performance. When applying our model to piano to multi-track arrangement, $\gamma$ balances the force of a noisy factor added to the piano, which encourages a partial unconditional generation capability. The experimental settings are the same as Section 5.5. We evaluate the results based on faithfulness and DOA. In Table 6, we report mean value, standard error of mean ($\mathrm{sem}$), and statistical significance computed by Wilcoxon signed rank test. By varying the $\gamma$ value, we observe a controllable balance between faithfulness and creativity. Specifically, a larger $\gamma$ encourages creativity (higher DOA) at the cost of faithfulness. If not mentioned otherwise, we use $\gamma=0.25$ for experiments in this paper.
@@ Appendix G Broader Impacts
However, we acknowledge the need to address potential risks. The accessibility of our system may inadvertently lead to excessive reliance on automation, potentially impeding the development of fundamental skills among musicians. Additionally, widespread adoption of the system may contribute to the homogenization of music, threatening the distinctiveness and individuality that are crucial to artistic expression. We recognize that our datasets predominantly feature contemporary Western music, which introduces a cultural bias that could limit the diversity of generated compositions.